# AthletePose3D: A Benchmark Dataset for 3D Human Pose Estimation and Kinematic Validation in Athletic Movements

### v2

@@ Abstract
Human pose estimation is a critical task in computer vision and sports biomechanics, with applications spanning sports science, rehabilitation, and biomechanical research. While significant progress has been made in monocular 3D pose estimation, current datasets often fail to capture the complex, high-acceleration movements typical of competitive sports. In this work, we introduce AthletePose3D, a novel dataset designed to address this gap. AthletePose3D includes 12 types of sports motions across various disciplines, with approximately 1.3 million frames and 165 thousand individual postures, specifically capturing high-speed, high-acceleration athletic movements. We evaluate state-of-the-art (SOTA) monocular 2D and 3D pose estimation models on the dataset, revealing that models trained on conventional datasets perform poorly on athletic motions. However, fine-tuning these models on AthletePose3D notably reduces the SOTA model mean per joint position error (MPJPE) from 214mm to 65mm—a reduction of over 69%. We also validate the kinematic accuracy of monocular pose estimations through waveform analysis, highlighting strong correlations in joint angle estimations but limitations in velocity estimation. Our work provides a comprehensive evaluation of monocular pose estimation models in the context of sports, contributing valuable insights for advancing monocular pose estimation techniques in high-performance sports environments. The dataset, code, and model checkpoints are available at: https://github.com/calvinyeungck/AthletePose3D
@@ 2 Related work
However, existing datasets are often not specifically designed for monocular 3D pose estimation or lack high-speed, high-acceleration movements that are essential in competitive sports. To address these limitations, we introduce AthletePose3D, a comprehensive dataset encompassing 12 types of sports actions performed by athletes ranging from amateurs (athletes without remuneration) to professionals. See Section 4 for more details.
@@ 2.2 Monocular 2D & 3D human pose estimations
In 2D pose estimation models, early methods were primarily CNN-based, proposing architectures such as ResNet [74], HRNet [60], and SwinPose [39]. Later approaches introduced transformers [69] to refine feature decoding from CNN backbones, including cascade transformers [28], TransPose [77], TokenPose [35], and UniFormer [29]. Further advancements replaced CNNs entirely with transformer-based architectures, as seen in HRFormer [81] and ViTPose [75], which employ vision transformers [12] for pose estimation. More recently, the resurgence of CNN models, such as ConvNets [40, 38, 53, 76], has led to a renewed focus on CNN-based methods like RTMPose [23] and MogaNet [32].
@@ 3.3 Estimated pose kinematic validation
The best-performing monocular 2D and 3D pose estimation models on the AthletePose3D validation set (see Sections 5.1 and 5.2) were selected for evaluation. Using these models, 2D and 3D poses were estimated from clips in the AthletePose3D test set. Next, limb joint velocity and joint angles were computed, as these joints are central to kinematic analysis and tend to have the highest errors. To smooth the kinematic data, a fourth-order Butterworth low-pass digital filter with zero-phase shift at 8 Hz was applied, following [62, 42].
Since coordination differences affect absolute kinematic values, the analysis focused on the waveform of kinematic data (e.g., the temporal series of wrist velocity in a javelin throw) rather than exact values, as the waveform can reveal valuable insights. Moreover, monocular pose estimation provides a more efficient alternative to conventional motion capture methods. To validate the motion waveform, paired t-test statistical parametric mapping (SPM) with a significance level of $\alpha=0.05$ was performed, and Pearson correlation was reported for a more intuitive interpretation.
@@ 4.2 Compare with existing dataset
Another key distinction between 3D human pose datasets, often overlooked, is the video length, as shown in Figure 4. Datasets such as Human3.6M and MPI-INF-3DHP, which focus on everyday movements, typically feature videos that extend beyond a thousand frames. In contrast, datasets capturing sports motions, like SportPose and AthletePose3D, usually consist of videos with around 250 frames. This is because sports motions are performed quickly, often within a matter of seconds, and each video typically captures only the specific action of interest. Given this, the use of 243 frames of data as input for traditional monocular 3D pose estimation models may not be the most suitable configuration. Therefore, we utilize the next supported and suitable input size, 81 frames.
@@ 5 Experiments results
MotionAGFormer H3.6M 223.47 197.46 235.68 62.09 302.74 535.56 237.43 91.04
(WACV’2024) [44] AP3D 76.95 73.82 58.49 17.00 79.06 122.64 67.23 10.00
H3.6M+AP3D 75.90 72.47 55.60 15.47 76.67 119.35 65.55 8.37
TCPFormer H3.6M 232.54 222.20 282.86 67.05 234.81 293.32 213.64 93.82
(AAAI’2025) [37] AP3D 76.01 72.11 55.95 16.09 77.07 119.52 65.80 8.15
H3.6M+AP3D 75.79 72.23 54.45 15.66 76.04 117.28 65.07 7.68
Table 4: Monocular 3D pose estimation performance on AthletePose3D validation set, with the top-performing result highlighted in bold. Each column for body parts indicates the average JPE for the specified body parts, while the MPJPE and P-MPJPE are also known as Protocols 1 and 2, respectively.
@@ 5.2 Monocular 3D pose estimation models performance
The results in Table 4 demonstrated the impact of different training datasets on monocular 3D pose estimation performance when evaluated on the AthletePose3D (AP3D) validation set. Models trained solely on Human3.6M (H3.6M) performed significantly worse than those trained on AP3D or a combination of both datasets. MotionAGFormer, when trained on H3.6M, exhibited high joint position errors, particularly for the wrist (235.68 mm) and ankle (535.56 mm), resulting in an MPJPE of 237.43 mm. Similarly, TCPFormer trained on H3.6M struggled with high errors, especially in the wrist (282.86 mm), although it achieved a slightly lower MPJPE of 213.64 mm.
Training on AP3D substantially improved performance for both models, reducing MPJPE by over 70% compared to training solely on H3.6M. TCPFormer achieved an MPJPE of 65.80 mm and a P-MPJPE of 8.15 mm, outperforming MotionAGFormer (67.23 mm MPJPE, 10.00 mm P-MPJPE). The most notable improvements were observed in lower-body joints, particularly the knee and ankle, where errors dropped from over 300 mm (H3.6M-trained models) to approximately 76 mm and 119 mm, respectively.
Utilizing both H3.6M and AP3D for training led to further refinements, with TCPFormer achieving the lowest overall errors (MPJPE: 65.07 mm, P-MPJPE: 7.68 mm). The best per-joint accuracy was also observed in this setting, with notable improvements in the wrist (54.45 mm), knee (76.04 mm), and ankle (117.28 mm). These results highlighted that training on sports-specific datasets like AP3D was crucial for achieving accurate monocular pose estimation in athletic movements, and combining diverse datasets further enhanced robustness.
@@ 5.3 Estimated pose kinematic validation
Upper limbs 0.47 0.032 0.29 0.021
Lower limbs 0.45 0.027 0.12 0.004
@@
Upper limbs 0.75 0.021 0.89 0.013
Lower limbs 0.66 $<$0.001 0.81 $<$0.001
@@
For velocity estimation, the 2D pose model showed moderate correlation with GT data for both the upper ($r=0.47$, $p=0.032$) and lower limbs ($r=0.45$, $p=0.027$), despite significant differences detected by the paired t-test SPM. Similarly, the 3D pose model exhibited a lower correlation for upper limb velocity ($r=0.29$, $p=0.021$) and a weak correlation for lower limb velocity ($r=0.12$, $p=0.004$), with significant waveform differences. These results suggested that while the estimated velocity waveforms captured some temporal trends, they still significantly deviated from GT motion capture.
For joint angles, the correlations were substantially higher, particularly for the 3D pose model. The 2D model achieved strong correlations for upper limb ($r=0.75$, $p=0.021$) and lower limb ($r=0.66$, $p<0.001$) joint angles. The 3D model further improved upon this, showing very strong correlations for upper limb ($r=0.89$, $p=0.013$) and lower limb ($r=0.81$, $p<0.001$) joint angles. However, despite these high correlations, all joint angle waveforms still exhibited statistically significant differences from GT data according to the paired t-test SPM, indicating residual discrepancies in waveform shape.

### v3

@@ Abstract
Human pose estimation is a critical task in computer vision and sports biomechanics, with applications spanning sports science, rehabilitation, and biomechanical research. While significant progress has been made in monocular 3D pose estimation, current datasets often fail to capture the complex, high-acceleration movements typical of competitive sports. In this work, we introduce AthletePose3D, a novel dataset designed to address this gap. AthletePose3D includes 12 types of sports motions across various disciplines, with approximately 1.3 million frames and 165 thousand individual postures, specifically capturing high-speed, high-acceleration athletic movements. We evaluate state-of-the-art (SOTA) monocular 2D and 3D pose estimation models on the dataset, revealing that models trained on conventional datasets perform poorly on athletic motions. However, fine-tuning these models on AthletePose3D notably reduces the SOTA model mean per joint position error (MPJPE) from 234mm to 98mm—a reduction of around 60%. We also validate the kinematic accuracy of monocular pose estimations through waveform analysis, highlighting strong correlations in joint angle estimations but limitations in velocity estimation. Our work provides a comprehensive evaluation of monocular pose estimation models in the context of sports, contributing valuable insights for advancing monocular pose estimation techniques in high-performance sports environments. The dataset, code, and model checkpoints are available at: https://github.com/calvinyeungck/AthletePose3D
@@ 2 Related work
However, existing datasets are often not specifically designed for monocular 3D pose estimation or lack high-speed, high-acceleration movements that are essential in competitive sports. To address these limitations, we introduce AthletePose3D, a comprehensive dataset encompassing 12 types of sports actions performed by athletes ranging from amateur athletes (athletes without remuneration) to professional athletes. See Section 4 for more details.
@@ 2.2 Monocular 2D & 3D human pose estimations
In 2D pose estimation models, early methods were primarily CNN-based, proposing architectures such as ResNet [74], HRNet [60], and SwinPose [39]. Later approaches introduced transformers [69] to refine feature decoding from CNN backbones, including cascade transformers [28], TransPose [77], TokenPose [35], and UniFormer [29]. Further advancements replaced CNNs entirely with transformer-based architectures, as seen in HRFormer [81] and ViTPose [75], which employ vision transformers [12] for pose estimation. More recently, the resurgence of CNN models [40, 38, 53, 76], has led to a renewed focus on CNN-based methods like RTMPose [23] and MogaNet [32].
@@ 3.3 Estimated pose kinematic validation
In general, monocular pose estimation provides a more efficient alternative to conventional motion capture methods. Therefore, the best-performing monocular 2D and 3D pose estimation models on the AthletePose3D validation set (see Sections 5.1 and 5.2) were selected for evaluation. Using these models, 2D and 3D poses were estimated from clips in the AthletePose3D test set. Next, limb joint velocity and joint angles were computed, as these joints are central to kinematic analysis and tend to have the highest errors. To smooth the kinematic data, a fourth-order Butterworth low-pass digital filter with zero-phase shift at 8 Hz was applied, following [62, 42].
Since coordination differences affect absolute kinematic values, the analysis focused on the waveform of kinematic data (e.g., the temporal series of wrist velocity in a javelin throw) rather than exact values, as the waveform can reveal valuable insights. To validate the motion waveform, paired t-test statistical parametric mapping (SPM) with a significance level of $\alpha=0.05$ was performed, and Pearson correlation was reported for a more intuitive interpretation.
@@ 4.2 Compare with existing dataset
Another key distinction between 3D human pose datasets, often overlooked, is the video length, as shown in Figure 4. Datasets such as Human3.6M and MPI-INF-3DHP, which focus on everyday movements, typically feature videos that extend beyond a thousand frames. In contrast, datasets capturing sports motions, like SportPose and AthletePose3D, generally consist of shorter videos, averaging around 250 frames. This reflects the dynamic nature of sports movements, which occur within seconds and are designed to capture specific actions of interest. Given this characteristic, we selected a shorter input size for the 3D pose estimation models utilized in this study, 81 frames.
@@ 5 Experiments results
MotionAGFormer H3.6M 250.09 234.27 254.57 74.87 320.39 560.23 257.26 95.67
(WACV’2024) [44] AP3D 118.58 122.45 88.48 30.45 112.84 168.10 100.12 31.84
H3.6M+AP3D 117.59 121.82 86.83 29.63 110.08 165.16 98.62 30.54
TCPFormer H3.6M 260.41 257.32 299.24 80.47 253.15 324.96 234.20 101.80
(AAAI’2025) [37] AP3D 117.16 121.25 87.56 30.24 110.40 165.15 98.77 30.67
H3.6M+AP3D 117.24 120.87 86.08 29.39 110.13 164.29 98.26 29.91
Table 4: Monocular 3D pose estimation (with 2D pose ground truth) performance on AthletePose3D validation set, with the top-performing result highlighted in bold. Each column for body parts indicates the average JPE for the specified body parts, while the MPJPE and P-MPJPE are also known as Protocols 1 and 2, respectively.
@@ 5.2 Monocular 3D pose estimation models performance
The results in Table 4 demonstrated the impact of different training datasets on monocular 3D pose estimation performance when evaluated on the AthletePose3D (AP3D) validation set. Models trained solely on Human3.6M (H3.6M) performed significantly worse than those trained on AP3D or a combination of both datasets. MotionAGFormer, when trained on H3.6M, exhibited high joint position errors, particularly for the Knee (320.39 mm) and ankle (560.23 mm), resulting in an MPJPE of 257.26 mm. Similarly, TCPFormer trained on H3.6M struggled with high errors, especially in the Ankle (324.96 mm), although it achieved a slightly lower MPJPE of 234.20 mm.
Training on AP3D substantially improved performance for both models, reducing MPJPE by around 60% compared to training solely on H3.6M. TCPFormer achieved an MPJPE of 98.77 mm and a P-MPJPE of 30.67 mm, outperforming MotionAGFormer (100.12 mm MPJPE, 31.84 mm P-MPJPE). The most notable improvements were observed in lower-body joints, particularly the knee and ankle, where errors dropped from over 300 mm (H3.6M-trained models) to approximately 110 mm and 170 mm, respectively.
Utilizing both H3.6M and AP3D for training led to further refinements, with TCPFormer achieving the lowest overall errors (MPJPE: 98.26 mm, P-MPJPE: 29.91 mm). The best per-joint accuracy was also observed in this setting, with notable improvements in the wrist (86.08 mm), knee (110.13 mm), and ankle (164.29 mm). These results highlighted that training on sports-specific datasets like AP3D was crucial for achieving accurate monocular pose estimation in athletic movements, and combining diverse datasets further enhanced robustness.
@@ 5.3 Estimated pose kinematic validation
Upper limbs 0.47 0.032 0.28 0.036
Lower limbs 0.45 0.027 0.11 0.004
@@
Upper limbs 0.75 0.021 0.90 0.001
Lower limbs 0.66 $<$0.001 0.82 0.044
@@
For velocity estimation, the 2D pose model showed moderate correlation with GT data for both the upper ($r=0.47$, $p=0.032$) and lower limbs ($r=0.45$, $p=0.027$), despite significant differences detected by the paired t-test SPM. Similarly, the 3D pose model exhibited a lower correlation for upper limb velocity ($r=0.28$, $p=0.036$) and a weak correlation for lower limb velocity ($r=0.11$, $p=0.004$), with significant waveform differences. These results suggested that while the estimated velocity waveforms captured some temporal trends, they still significantly deviated from GT motion capture.
For joint angles, the correlations were substantially higher, particularly for the 3D pose model. The 2D model achieved strong correlations for upper limb ($r=0.75$, $p=0.021$) and lower limb ($r=0.66$, $p<0.001$) joint angles. The 3D model further improved upon this, showing very strong correlations for upper limb ($r=0.90$, $p=0.001$) and lower limb ($r=0.82$, $p=0.044$) joint angles. However, despite these high correlations, all joint angle waveforms still exhibited statistically significant differences from GT data according to the paired t-test SPM, indicating residual discrepancies in waveform shape.